OpenStudy (anonymous):

What is the meaning to find the 15th term in the geometric sequence.

OpenStudy (ranga):

What is the formula for the nth term of a geometric sequence?

OpenStudy (ranga):

\[\Large a_{n} = a_{1}r^{n-1}\quad a _{1} = \text{First Term;} \text{ r = common ratio}\]

OpenStudy (ranga):

Here the first term = 20 common ratio r = 10/20 = 0.5 n = 15 Plug it into the formula and you will have your 15th term.